# Discrete Approximate Circle Bundles

### v2

@@ 3.1 Three Characterizations
Given a paracompact space $B$ and $\varepsilon\geq 0$, an $\varepsilon$-approximate classifying map is a map $f:B\to\text{Gr}(2)^{\varepsilon}$.
@@ 3.2 Relationships Between Notions
$\text{wit}:T_{\alpha}(\pi,\mathcal{U})\to\check{Z}^{1}_{(3\sqrt{2})\alpha}(\mathcal{U};\underline{O(2)})\hskip 28.45274pt\text{cl}:\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})\to\text{Maps}(B,\text{Gr}(2)^{\sqrt{2}\varepsilon})\\$
We first show that discrete approximate local trivializations are witnessed by discrete approximate cocycles:
@@
Suppose $\Omega$ witnesses that $\{f_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$ for some $\alpha\geq 0$. Then $\Omega\in\check{Z}^{1}_{3\alpha\sqrt{2}}(\mathcal{U};\underline{O(2)})$.
@@
Assume $\{f_{j}\}\in T_{\varepsilon,\delta}(\pi,\mathcal{U})$, where $\alpha=\frac{\varepsilon}{1-\delta}$. Suppose $x\in\pi^{-1}(U_{j}\cap U_{k}\cap U_{l})$. Let $v_{\tau\tau^{\prime}}=\Omega_{\tau\tau^{\prime}}f_{\tau^{\prime}}(x)-f_{\tau}(x)$ for $\tau,\tau^{\prime}\in\{j,k,l\}$. Then,
$\left|\Omega_{jk}\Omega_{kl}f_{l}(x)-\Omega_{jl}f_{l}(x)\right|=\left|\Omega_{jk}(v_{kl}+f_{k}(x))-\Omega_{jl}f_{l}(x)\right|$
$\left|\Omega_{jk}v_{kl}+(f_{j}(x)+v_{jk})-(f_{j}(x)-v_{jl})\right|=\left|\Omega_{jk}v_{kl}+v_{jk}-v_{jl}\right|$
$\leq|v_{kl}|+|v_{jk}|+|v_{jl}|<3\varepsilon\\$
The result then follows immediately from Proposition A.10.
Next, we show how to algorithmically compute a witness for a discrete approximate circle bundle.
Construction 3.5.
Let $X=\{x_{s}\}_{s=1}^{n}$ be a finite set, and let $f,g:S\to\mathbb{S}^{1}$. Compute the values $\{\alpha_{s}=\cos^{-1}(f(x_{s}))\}^{n}_{s=1}$ and $\{\beta_{s}=\cos^{-1}(g(x_{s}))\}^{n}_{s=1}$, then define $\gamma_{s}=\alpha_{s}-\beta_{s}$ and $\tilde{\gamma}_{s}=\alpha_{s}+\beta_{s}$ for all $s$. Find the respective midpoints $\theta$ and $\widetilde{\theta}$ of the shortest arc $C$ containing $\{\gamma_{s}\}_{s=1}^{n}$ and the shortest arc $\widetilde{C}$ containing $\{\widetilde{\gamma}_{s}\}_{s=1}^{n}$. Compute the values
$d_{\text{rot}}=\text{max}_{s}|f(x_{s})-R_{\theta}g(x_{s})|,\hskip 28.45274ptd_{\text{ref}}=\text{max}_{s}|f(x_{s})-R_{\widetilde{\theta}}g(x_{s})|$
If $d_{\text{rot}}<d_{\text{ref}}$, set $\Omega=R_{\theta}$; otherwise, set $\Omega=R_{\widetilde{\theta}}r$, where $R_{\theta}=\left(\begin{array}[]{cc}\cos(\theta)&-\sin(\theta)\\ \sin(\theta)&\cos(\theta)\\ \end{array}\right)$ and $r=\left(\begin{array}[]{cc}1&0\\ 0&-1\\ \end{array}\right)$.
@@
If $n>1$ and $\text{diam}(\{\gamma_{s}\}_{s=1}^{n})<\pi$ in Construction 3.5 above, then $\Omega\in O(2)$ is the unique minimizer of $\text{max}_{s}|f(x_{s})-\Omega g(x_{s})|$.
@@
Let $L:O(2)\to\mathbb{R}_{\geq 0}$ be defined by $L(\Omega)=\max_{s}|f(x_{s})-\Omega g(x_{s})|^{2}$. We have
$|f(x_{s})-\Omega g(x_{s})|^{2}$ $=|f(x_{s})|^{2}+|g(x_{2})|^{2}-2f(x_{s})\cdot\Omega g(x_{s})$
$=2-2f(x_{s})\cdot\Omega g(x_{s})$
for all $s$. If $\Omega=R_{\theta}$ is a counter-clockwise rotation by $\theta$, then $f(x_{s})\cdot\Omega g(x_{s})=\cos(\gamma_{s}-\theta)$, so to minimize $L$ over $SO(2)$ we need to identify $\theta\in\mathbb{S}^{1}$ for which the angular distance to the farthest $\gamma_{s}$ is minimized. If the measure of $C$ is less than $\pi$, then this is precisely achieved at the midpoint of the shortest arc $C$ containing $\{\gamma_{s}\}_{s=1}^{n}$ (furthermore, the requirement that $\text{diam}(\{\gamma_{s}\}_{s=1}^{n})<\pi$ guarantees that $C$ is unique). To see this, first note that the optimal $\theta$ certainly lies in $C$, and for fixed $\theta\in C$, $\cos(\gamma_{s}-\theta)$ is smallest when $\gamma_{s}$ is an endpoint of $C$. By symmetry, the $\theta$ which maximizes the minimum of the values $\{\cos(\gamma_{s}-\theta)\}_{s=1}^{n}$ must be half way between the two end points (otherwise, a small perturbation of $\theta$ towards the midpoint would move $\theta$ closer to the more distant endpoint while still keeping $\theta$ nearer to the closer endpoint. This would yield a more optimal choice for $\theta$).
Analogously, if $\Omega=R_{\widetilde{\theta}}r$ for some $\widetilde{\theta}\in\mathbb{R}$, one can check that $f(x_{s})\cdot\Omega g(x_{s})=\cos(\widetilde{\gamma}_{s}-\theta)$. Thus, to minimize $L$ over the component of $O(2)$ with matrices of negative determinant, we need to identify $\widetilde{\theta}\in\mathbb{S}^{1}$ for which the angular distance to the farthest $\widetilde{\gamma}_{s}$ is minimized. Finally, comparing $d_{\text{rot}}$ and $d_{\text{ref}}$ determines which component of $O(2)$ contains the optimal $\Omega$. To see that these two values cannot be equal, observe that if $\text{diam}(\{\gamma_{s}\}_{s=1}^{n})<\pi$, then $\text{diam}(\{\widetilde{\gamma}_{s}\}_{s=1}^{n})>\pi$. It follows that then $d_{\text{rot}}<\frac{\pi}{2}$ and $d_{\text{ref}}>\frac{\pi}{2}$. Similarly, if $\text{diam}(\{\widetilde{\gamma}_{s}\}_{s=1}^{n})<\pi$, then $\text{diam}(\{\gamma_{s}\}_{s=1}^{n})>\pi$, so $d_{\text{ref}}<d_{\text{rot}}$.
@@
If $\{f_{j}\}\in T_{\varepsilon,\delta}(\pi,\mathcal{U})$ for $\varepsilon<\sqrt{2}$ and $|\pi^{-1}(U_{jk})|>1$ for all $(jk)\in\mathcal{N}(\mathcal{U})$, then Construction 3.5 yields a witness.
@@
$\gamma^{jk}_{s}=(f_{j}(x_{s}))\cdot(\Omega_{jk}f_{k}(x_{s}))>1-\frac{\varepsilon^{2}}{2}$
for all $x_{s}\in U_{j}\cap U_{k}\in\mathcal{N}(\mathcal{U})$. Then $|\gamma^{jk}_{s}|>\cos^{-1}(1-\frac{\varepsilon^{2}}{2})$ for all $x_{s}\in U_{j}\cap U_{k}\in\mathcal{N}(\mathcal{U})$, so $\text{diam}(\{\gamma^{jk}_{s}\})<2\cos^{-1}(1-\frac{\varepsilon^{2}}{2})$. We conclude that if $\cos^{-1}(1-\frac{\varepsilon^{2}}{2})<\frac{\pi}{2}$ (or, equivalently, if $\varepsilon<\sqrt{2}$), then each $\Omega_{jk}$ obtained using Construction 3.5 is the unique minimizer of $\text{max}_{x_{s}\in\pi^{-1}(U_{j}\cap U_{k})}|f_{j}(x_{s})-f_{k}(x_{s})|$. Thus by Proposition 3.4, $\Omega$ is a witness that $\{f_{j}\}\in T_{\varepsilon,\delta}(\pi,\mathcal{U})$.
@@
Suppose $\Omega,\Lambda\in\check{C}^{1}(\mathcal{U};\underline{O(2)})$ respectively witness $\{f_{j}\}\in T_{\varepsilon_{f},\infty}(\pi,\mathcal{U})$ and $\{g_{j}\}\in T_{\varepsilon_{g},\infty}(\pi,\mathcal{U})$, and $d_{h}(f_{k}(\pi^{-1}(U_{jk})),\mathbb{S}^{1})<\delta_{1}$ (with respect to the Euclidean metric) for every ordered 1-simplex $(jk)\in\mathcal{N}(\mathcal{U})$. If $d_{T}(\{f_{j}\},\{g_{j}\})<\delta_{2}$, then $d_{Z}(\Omega,\Lambda)<\frac{\sqrt{2}}{1-\delta_{1}}(\varepsilon_{f}+\varepsilon_{g}+2\delta_{2})$.
@@
Given any $(jk)\in\mathcal{N}(\mathcal{U})$ and $x\in\pi^{-1}(U_{j}\cap U_{k})$, the triangle inequality implies
$|\Omega_{jk}f_{k}(x)-\Lambda_{jk}f_{k}(x)|\leq$
@@
If $\Omega,\Lambda\in\check{C}^{1}(\mathcal{U};\underline{O(2)})$ both witness $\{f_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$ then $d_{Z}(\Omega,\Lambda)<(2\sqrt{2})\alpha$.
@@
Suppose $\Omega,\Lambda\in\check{C}^{1}(\mathcal{U};\underline{O(2)})$ respectively witness $\{f_{j}\},\{g_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$. If $d_{T}(\{f_{j}\},\{g_{j}\})<\delta$, then $d_{Z}(\Omega,\Lambda)<2\sqrt{2}(\alpha+\delta)$.
The following construction and subsequent result are due to [SP23]:
Construction 3.11.
Suppose $\Omega\in\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})$ for some locally-finite open cover $\mathcal{U}=\{U_{j}\}_{j\in J}$ of a paracompact space $B$. Choose a partition of unity $\{\rho_{j}\}_{j\in J}$ subordinate to $\mathcal{U}$. For each $j$, define $\Phi_{j}:U_{j}\to V(2)$ by
$\Phi_{j}(b)=\left(\begin{array}[]{ccc}\sqrt{\rho_{0}(b)}\Omega_{0j}(b)\\ \sqrt{\rho_{1}(b)}\Omega_{1j}(b)\\ \vdots\\ \end{array}\right)$ (15)
@@
$\widetilde{f}(b)=\sum_{j\in J}\rho_{j}(b)\Phi_{j}(b)\Phi_{j}(b)^{T}$ (16)
Since each $b$ is contained in a finite number of sets in $\mathcal{U}$, this map is well-defined.
For a proof of the proposition below, see [SP23]:
@@
For $\varepsilon<\frac{1}{2}$ and any partition of unity $\{\rho_{j}\}_{j\in J}$, Construction 3.11 yields a map $\text{cl}:\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})\to\text{Maps}(B,\text{Gr}(d)^{\sqrt{2}\varepsilon})$ with the following properties:
@@
    The induced map $\text{cl}:\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})\to[B,\text{Gr}(d)^{\sqrt{2}\varepsilon}]$ factors through $\check{H}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})$ and does not depend on choice of $\{\rho_{j}\}_{j\in J}$.
@@
    If $\Omega,\Lambda\in\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})$ such that $d_{Z}(\Omega,\Lambda)<\delta$, then $\text{cl}(\Omega)$ and $\text{cl}(\Lambda)$ become equal in $B,\text{Gr}(d)^{\sqrt{2}(\delta+\varepsilon)}]$.

### v3

@@ 3.3 Approximate Classifying Maps
Given a space $B$ and $\varepsilon\geq 0$, an $\varepsilon$-approximate classifying map is a map $f:B\to\text{Gr}(2)^{\varepsilon}$, where $\text{Gr}(2)^{0}=\text{Gr}(2)$, and $\text{Gr}(2)^{\varepsilon}\subseteq\mathbb{R}^{\infty\times\infty}$ for $\varepsilon>0$ is the $\varepsilon$-thickening
$\text{Gr}(2)^{\varepsilon}=\{A\in\mathbb{R}^{\infty\times\infty}:\|A-P\|_{F}<\varepsilon\text{ for some }P\in\text{Gr}(2)\}\\$
@@ 3.4 Relationships Between Notions
$\text{wit}:\Gamma_{\alpha}(\pi,\mathcal{U})\to\check{Z}^{1}_{(3\sqrt{2})\alpha}\left(\mathcal{U};\underline{O(2)}\right)\hskip 28.45274pt\text{cl}:\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})\to\text{Maps}\left(B,\text{Gr}(2)^{\sqrt{2}\varepsilon}\right)\\$
We first show that systems of discrete approximate local angle functions are witnessed by discrete approximate cocycles:
@@
Suppose $\Omega$ witnesses that $\mathfrak{F}\in\Gamma_{\alpha}(\pi,\mathcal{U})$ for some $\alpha>0$. Then $\Omega\in\check{Z}^{1}_{3\alpha\sqrt{2}}\left(\mathcal{U};\underline{O(2)}\right)$.
@@
Assume $\mathfrak{F}=\{f_{j}\}_{j\in J}\in\Gamma_{\varepsilon,\delta}(\pi,\mathcal{U})$, where $\alpha=\frac{\varepsilon}{1-\delta}$. Suppose $x\in\pi^{-1}(U_{jkl})$ and let $v_{st}=\Omega_{st}f_{t}(x)-f_{s}(x)$ for $s,t\in\{j,k,l\}$. Then,
$\left|\Omega_{jk}\Omega_{kl}f_{l}(x)-\Omega_{jl}f_{l}(x)\right|\quad=\quad\left|\Omega_{jk}(v_{kl}+f_{k}(x))-\Omega_{jl}f_{l}(x)\right|$
$\quad\left|\Omega_{jk}v_{kl}+(f_{j}(x)+v_{jk})-(f_{j}(x)-v_{jl})\right|\quad=\quad\left|\Omega_{jk}v_{kl}+v_{jk}-v_{jl}\right|$
$\leq\quad|v_{kl}|+|v_{jk}|+|v_{jl}|\quad<\quad 3\varepsilon\\$
The result then follows from applying Proposition A.12 to each matrix $(\Omega_{jk}\Omega_{kl}-\Omega_{jl})$.
Below we describe an explicit construction which produces a witness for a system of discrete approximate local angle functions.
Construction 3.16.
Let $X$ be any nonempty set, and let $f,g:X\to\mathbb{S}^{1}$. For each $x\in X$, define $e^{i\alpha_{x}}=f(x)$ and $e^{i\beta_{x}}=g(x)$, then define $\gamma_{x}=\alpha_{x}-\beta_{x}$ and $\tilde{\gamma}_{x}=\alpha_{x}+\beta_{x}$. Let $e^{i\theta}$ denote the midpoint of the shortest arc $C$ in $\mathbb{S}^{1}$ containing $\{e^{i\gamma_{x}}\}_{x\in X}$, and let $e^{i\widetilde{\theta}}$ denote the midpoint of the shortest arc $\widetilde{C}$ containing $\{e^{i\widetilde{\gamma}_{x}}\}_{x\in X}$. Define
$d_{\text{rot}}=\sup_{x\in X}|f(x)-R_{\theta}g(x)|,\hskip 28.45274ptd_{\text{ref}}=\sup_{x\in X}|f(x)-R_{\widetilde{\theta}}g(x)|$
If $d_{\text{rot}}<d_{\text{ref}}$, set $\Omega^{*}=R_{\theta}$; otherwise, set $\Omega^{*}=R_{\widetilde{\theta}}r$, where $R_{\theta}=\left(\begin{array}[]{cc}\cos(\theta)&-\sin(\theta)\\ \sin(\theta)&\cos(\theta)\\ \end{array}\right)$ and $r=\left(\begin{array}[]{cc}1&0\\ 0&-1\\ \end{array}\right)$.
@@
[…]
@@
Let $L:O(2)\to\mathbb{R}_{\geq 0}$ be defined by $L(\Omega)=\sup\limits_{x\in X}|f(x)-\Omega g(x)|^{2}$. We have
$|f(x)-\Omega g(x)|^{2}\quad$ $=\quad|f(x)|^{2}+|g(x)|^{2}-2f(x)\cdot\Omega g(x)$
$=\quad 2-2f(x)\cdot\Omega g(x)$
for all $x$ (here we are identifying $\mathbb{R}^{2}$ with $\mathbb{C}$ in the usual way). If $\Omega=R_{\theta}$ is a counter-clockwise rotation by $\theta$, then $f(x)\cdot\Omega g(x)=\cos(\gamma_{x}-\theta)$, so to minimize $L$ over $SO(2)$ we need to identify $\theta\in\mathbb{S}^{1}$ for which the angular distance to the farthest $\gamma_{x}$ is minimized. If the measure of $C$ is less than $\pi$, then this is precisely achieved at the midpoint of the shortest arc $C$ containing $\{e^{i\gamma_{x}}\}_{x\in X}$ (furthermore, the requirement that $\text{diam}(\{e^{i\gamma_{x}}\}_{x\in X})<\pi$ guarantees that $C$ is unique). To see this, first note that $e^{i\theta}$ certainly lies in $C$ for any optimal $\theta$, and for any fixed $e^{i\theta}\in C$, $d_{\mathbb{S}^{1}}(e^{i\gamma_{x}},e^{i\theta})$ is smallest when $e^{i\gamma_{x}}$ is an endpoint of $C$. By symmetry, the $\theta$ which maximizes the minimum of the values $\{d_{\mathbb{S}^{1}}(e^{i\gamma_{x}},e^{i\theta})\}_{x\in X}$ must be half way between the two end points (otherwise, a small perturbation of $\theta$ towards the midpoint would move $\theta$ closer to the more distant endpoint while still keeping $\theta$ nearer to the closer endpoint. This would yield a more optimal choice for $\theta$).
Analogously, if $\Omega=R_{\widetilde{\theta}}r$ for some $\widetilde{\theta}\in\mathbb{R}$, one can check that $f(x)\cdot\Omega g(x)=\cos(\widetilde{\gamma}_{x}-\theta)$. Thus, to minimize $L$ over the component of $O(2)$ with matrices of negative determinant, we need to identify $\widetilde{\theta}\in\mathbb{S}^{1}$ for which the angular distance to the farthest $\widetilde{\gamma}_{x}$ is minimized. Finally, comparing $d_{\text{rot}}$ and $d_{\text{ref}}$ determines which component of $O(2)$ contains the optimal $\Omega$. To see that these two values cannot be equal, observe that if $\text{diam}(\{e^{i\gamma_{x}}\}_{x\in X})<\pi$, then $\text{diam}(\{e^{i\widetilde{\gamma}_{x}}\}_{x\in X})>\pi$. It follows that then $d_{\text{rot}}<\frac{\pi}{2}$ and $d_{\text{ref}}>\frac{\pi}{2}$. Similarly, if $\text{diam}(\{e^{i\widetilde{\gamma}_{x}}\}_{x\in X})<\pi$, then $\text{diam}(\{e^{i\gamma_{x}}\}_{x\in X})>\pi$, so $d_{\text{ref}}<d_{\text{rot}}$.
@@
If $\mathfrak{F}\in T_{\varepsilon,\delta}(\pi,\mathcal{U})$ for $\varepsilon<\sqrt{2}$ and $|\pi^{-1}(U_{jk})|>1$ for all $(jk)\in\mathcal{N}(\mathcal{U})$, then Construction 3.16 yields a witness.
@@
$\gamma^{jk}_{x}=(f_{j}(x))\cdot(\Omega_{jk}f_{k}(x))>1-\frac{\varepsilon^{2}}{2}$
for all $x\in U_{jk}$ and $(jk)\in\mathcal{N}(\mathcal{U})$. Then $|\gamma^{jk}_{x}|>\cos^{-1}(1-\frac{\varepsilon^{2}}{2})$ for all $x\in U_{jk}$, so $\text{diam}(\{\gamma^{jk}_{x}\})<2\cos^{-1}(1-\frac{\varepsilon^{2}}{2})$. We conclude that if $\cos^{-1}(1-\frac{\varepsilon^{2}}{2})<\frac{\pi}{2}$ (or, equivalently, if $\varepsilon<\sqrt{2}$), then each $\Omega_{jk}$ obtained using Construction 3.16 is the unique minimizer of $\max\limits_{x\in\pi^{-1}(U_{jk})}|f_{j}(x)-f_{k}(x)|$. Thus by Proposition 3.15, $\Omega$ is a witness that $\mathfrak{F}\in T_{\varepsilon,\delta}(\pi,\mathcal{U})$.
@@
Let $\delta<1$. Suppose $\Omega,\Lambda\in\check{C}^{1}\left(\mathcal{U};\underline{O(2)}\right)$ witness $\{f_{j}\}_{j\in J}\in\Gamma_{\varepsilon_{f},\delta}(\pi,\mathcal{U})$ and $\{g_{j}\}_{j\in J}\in\Gamma_{\varepsilon_{g},\delta}(\pi,\mathcal{U})$, respectively, and $d_{H}(f_{k}(\pi^{-1}(U_{jk})),\mathbb{S}^{1})<\delta_{1}$ in $(\mathbb{C},|\cdot|)$ for every ordered 1-simplex $(jk)\in\mathcal{N}(\mathcal{U})$. If $d_{\Gamma}(\{f_{j}\}_{j\in J},\{g_{j}\}_{j\in J})<\delta_{2}$, then $d_{Z}(\Omega,\Lambda)<\frac{\sqrt{2}}{1-\delta_{1}}(\varepsilon_{f}+\varepsilon_{g}+2\delta_{2})$.
@@
Given any $(jk)\in\mathcal{N}(\mathcal{U})$ and $x\in\pi^{-1}(U_{jk})$, the triangle inequality implies
$|\Omega_{jk}f_{k}(x)-\Lambda_{jk}f_{k}(x)|\quad\leq$
@@
If $\Omega,\Lambda\in\check{C}^{1}\left(\mathcal{U};\underline{O(2)}\right)$ both witness $\mathfrak{F}\in\Gamma_{\alpha}(\pi,\mathcal{U})$, then $d_{Z}(\Omega,\Lambda)<2\sqrt{2}\,\alpha$.
@@
Suppose $\Omega,\Lambda\in\check{C}^{1}\left(\mathcal{U};\underline{O(2)}\right)$ witness $\mathfrak{F},\mathfrak{G}\in T_{\alpha}(\pi,\mathcal{U})$, respectively. If $d_{\Gamma}(\mathfrak{F},\mathfrak{G})<\delta$, then $d_{Z}(\Omega,\Lambda)<2\sqrt{2}(\alpha+\delta)$.
The following construction and subsequent result are due to [24, Construction 4.15]:
Construction 3.22.
Suppose $\Omega\in\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})$ for some locally-finite open cover $\mathcal{U}=\{U_{j}\}_{j\in\mathbb{N}}$ of a paracompact space $B$. Choose a partition of unity $\{\rho_{j}\}_{j\in\mathbb{N}}$ subordinate to $\mathcal{U}$. For each $j\in\mathbb{N}=\{0,1,\ldots\}$, define $\Psi_{j}:U_{j}\to V(2)$ by
$\Psi_{j}(b)=\left(\begin{array}[]{ccc}\sqrt{\rho_{0}(b)}\Omega_{0j}(b)\\ \sqrt{\rho_{1}(b)}\Omega_{1j}(b)\\ \vdots\\ \end{array}\right)$ (21)
@@
$\widetilde{f}(b)=\sum_{j=0}^{\infty}\rho_{j}(b)\Psi_{j}(b)\Psi_{j}(b)^{T}$ (22)
Since each $b$ is contained in a finite number of sets in $\mathcal{U}$, the sum above is finite and hence $\tilde{f}$ is well-defined.
For a proof of the theorem below, see [24, Lemmas 4.16, 4.17, 4.18 and Theorem 4.21]:
@@
For $\varepsilon\geq 0$ and any partition of unity $\{\rho_{j}\}_{j\in\mathbb{N}}$, Construction 3.22 yields a map $\text{cl}:\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})\to\text{Maps}(B,\text{Gr}(d)^{\sqrt{2}\varepsilon})$ with the following properties:
@@
    The induced map $\text{cl}:\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})\to\left[B,\text{Gr}(d)^{\sqrt{2}\varepsilon}\right]$ factors through $\check{H}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})$ and does not depend on the choice of $\{\rho_{j}\}_{j\in\mathbb{N}}$.
@@
    If $\Omega,\Lambda\in\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})$ satisfy $d_{Z}(\Omega,\Lambda)<\delta$, then $\text{cl}(\Omega)$ and $\text{cl}(\Lambda)$ become equal in $\left[B,\text{Gr}(d)^{\sqrt{2}(\delta+\varepsilon)}\right]$.